# FALCON: An ML Framework for Fully Automated Layout-Constrained Analog Circuit Design

### v1

@@ 2 Related Work
To the best of our knowledge, FALCON is the first framework to unify topology selection, parameter inference, and layout-aware optimization in a single end-to-end pipeline, validated at scale using industrial-grade Cadence simulations for mm-wave analog circuits.
@@ 3.1 Dataset Overview
For each topology, expert-designed schematics were implemented in Cadence Virtuoso, and key design parameters were manually identified based on their functional relevance. Parameter ranges were specified by domain experts and systematically swept using Cadence ADE XL, enabling parallelized Spectre simulations across the design space. For each configuration, performance metrics—such as gain, bandwidth, and oscillation frequency—were extracted and recorded. Each datapoint therefore includes the full parameter vector, the corresponding Cadence netlist, and the simulated performance metrics. The resulting dataset comprises over one million datapoints, capturing a wide range of circuit behaviors and design trade-offs across diverse topologies. This large-scale, high-fidelity dataset forms the foundation for training and evaluating our inverse design pipeline.
@@ 3.2 Graph-Based Circuit Representation
To support learning over such structured graphs, each edge is annotated with a rich set of attributes: (i) a categorical device type, specifying the component and connected terminal pair (e.g., NMOS drain–gate, resistor); (ii) numeric attributes, such as channel length or port resistance, fixed by the schematic; (iii) parametric attributes, defined symbolically in the netlist (e.g., W1, R3) and resolved numerically during preprocessing; (iv) one-hot categorical features, such as source type (DC, AC, or none); and (v) computational attributes, such as diffusion areas (Ad, As) derived from sizing. This rule-based graph construction generalizes across circuit families without task-specific customization. Graphs in the FALCON dataset range from 4–40 nodes and 7–70 edges, reflecting the variability of practical analog designs.
@@ 4 Stage 1: Performance-Driven Topology Selection
We assess classification performance using accuracy, balanced accuracy, macro precision, macro recall, macro F1, and micro F1 scores on the test set. As summarized in Table 1, the classifier achieves an overall accuracy of 99.57%, with macro F1 of 99.30% and balanced accuracy of 99.33%, demonstrating strong generalization across all 20 circuit topologies. Micro F1 (identical to accuracy in the multiclass setting) reaches 99.57%, while macro metrics—averaged equally across classes—highlight robustness to class imbalance. These trends are reinforced by the per-class accuracy plot in Figure 3(c), where most topologies reach 100% accuracy.
@@ 5 Stage 2: Generalizable Forward Modeling for Performance Prediction
where $m_{i}=1$ if the $i$-th metric is defined for the current sample, and zero otherwise.
Model Architecture and Training. Each circuit is represented as an undirected multi-edge graph with voltage nets as nodes and circuit components as edges. All circuit parameters—both fixed and sweepable—are assigned to edges, along with categorical device types and one-hot encoded indicators. For each edge $(u,v)$, these attributes are concatenated to form a unified feature vector $x_{uv}$. The feature set is consistent within each component type but varies across types (e.g., NMOS vs. inductor), reflecting the structure defined in Section 3.2.
@@
To account for component heterogeneity, we apply type-specific MLP encoders $\phi_{\text{enc}}^{(t)}$ to each edge feature vector, producing initial embeddings $e_{uv}^{(0)}=\phi_{\text{enc}}^{(t)}(x_{uv})$, where $t$ is the component type. These embeddings are updated via a 4-layer edge-centric message-passing GNN with shared weights. At each layer $\ell$, for each node $u$, we first compute the node hidden state using the edge embeddings of all neighbors of the node $u$, $\mathcal{N}(u)$. Then, for each edge $(u,v)$ in the circuit graph, we compute the edge embedding at the next layer $\ell+1$ is using the edge embedding $e_{uv}^{(\ell)}$ and the hidden node states forming the edge $(u,v)$ at the current layer $\ell$ as follows:
$h_{u}^{(\ell)}=\sum_{w\in\mathcal{N}(u)}\phi_{\text{MSG}}(e_{wu}^{(\ell)}),\quad e_{uv}^{(\ell+1)}=\phi_{\text{UPD}}(e_{uv}^{(\ell)},h_{u}^{(\ell)},h_{v}^{(\ell)}),$
where $\phi_{\text{MSG}},\phi_{\text{UPD}}$ are the message and update parameters of message-passing GNN and $h_{u}^{(\ell)},h_{v}^{(\ell)}$ are the hidden states for the nodes forming the edge $(u,v)$ respectively. After message passing [57], final edge embeddings $e_{uv}^{(L)}$ are aggregated to form a graph-level representation $z_{\text{graph}}=\sum_{(u,v)}e_{uv}^{(L)}$, which is decoded by a fully connected MLP (hidden size 256) to predict the 16-dimensional performance vector $\hat{y}\in\mathbb{R}^{16}$. An overview of this GNN-based forward prediction pipeline is shown in Figure 4.
@@
Evaluation. We evaluate the accuracy of the GNN forward model $f_{\theta}$ on a test set drawn from 19 of the 20 topologies. One topology—RVCO—is entirely excluded from training, validation, and test splits to assess generalization to unseen architectures. Prediction quality is measured using standard regression metrics: coefficient of determination ($\text{R}^{2}$), root mean squared error (RMSE), and mean absolute error (MAE), computed independently for each of the 16 performance metrics. We also report the mean relative error per metric, computed as the average across all test samples where each metric is defined. As summarized in Table 2, the model achieves high accuracy across all dimensions, with an average $\text{R}^{2}$ of 0.971.
To evaluate end-to-end prediction accuracy at the sample level, we compute the mean relative error per instance, defined as the average relative error across all valid (non-masked) performance metrics for each test sample. Figure 5 shows the distribution of this quantity across the test set (trimmed at the 95th percentile to reduce the impact of outliers). The distribution is sharply concentrated, indicating that most predictions closely match their corresponding target vectors. Without percentile trimming, the overall mean relative error across the full test set is 9.14%.
@@
1.0 1.0 0.99 1.0 0.92 1.0 0.99 0.98 0.97 0.83 0.97 1.0 1.0 1.0 0.89 1.0
RMSE 0.289 0.107 0.536 0.84 1.517 0.206 0.534 0.969 0.721 0.293 0.908 0.1 0.232 0.144 2.541 0.07
MAE 0.212 0.077 0.208 0.188 0.554 0.116 0.202 0.369 0.181 0.097 0.232 0.069 0.168 0.104 1.167 0.046
Rel. Err. 11.5% 2.7% 18.6% 7.8% 11.4% 1.8% 4.5% 5.6% 0.6% 6.5% 4.4% 4.5% 4.5% 11.7% 1.3% 1.48%
@@ 6 Stage 3: Layout-Aware Parameter Inference via Gradient Reasoning
where $\mathcal{L}_{\text{perf}}$ is the masked mean squared error (see Eqn 2) between predicted and target performance vectors, and $\mathcal{L}_{\text{layout}}$ is a normalized area penalty derived from analytical layout equations. To prioritize functionality, layout loss is softly gated by:
@@
which attenuates layout penalties when performance error exceeds a threshold $\tau$, encouraging the model to first achieve functionality before optimizing for layout compactness.
We set $\tau=0.05$, $\gamma=50$, and normalize layout area by $1\,\text{mm}^{2}$ to stabilize gradients. The layout weight $\lambda_{\text{area}}=0.02$ is chosen empirically to balance performance accuracy and physical realism without dominating the loss. This gated formulation supports manufacturable parameter recovery and reflects the broader paradigm of physics-informed learning [58].
@@
The layout model is deterministic and non-learned. It estimates area contributions from passive components—capacitors, inductors, and resistors—as these dominate total area and exhibit layout-sensitive behavior. Active devices (e.g., MOSFETs) are excluded since their geometries are fixed by the PDK and are negligible [63]. For a given parameter vector $x$, the total layout loss is computed as:
$\mathcal{L}_{\text{layout}}(x)=\sum_{e\in\mathcal{E}_{\text{passive}}}A_{e}(x),$
where $\mathcal{E}_{\text{passive}}$ is the set of passive elements, and $A_{e}(x)$ is the area of the created layout for the passive component based on analytical physics-based equations. The area of element $e$ is estimated based on its 2D dimensions (e.g., $A=W\cdot L$ for resistors and capacitors). This area is normalized and used as a differentiable penalty in the optimization objective (see Eqn 3).
Gradient Reasoning Procedure. Starting from the initialized parameter vector $x$, we iteratively update parameters via gradient reasoning. At each step, the frozen forward model $f_{\theta}$ predicts the performance $\hat{y}=f_{\theta}(T,x)$, and the total loss $\mathcal{L}_{\text{total}}$ is evaluated. Gradients are backpropagated with respect to $x$, and updates are applied using the Adam optimizer. Optimization proceeds for a fixed number of steps, with early stopping triggered if the loss fails to improve over a predefined window.
@@ 7 Conclusion and Future Work
We presented FALCON, a modular framework for end-to-end analog and RF circuit design that unifies topology selection, performance prediction, and layout-aware parameter optimization. Trained on over one million Cadence-simulated mm-wave circuits, FALCON combines a lightweight MLP, a generalizable GNN, and differentiable gradient reasoning to synthesize circuits from specification to layout-constrained parameters. FALCON achieves >99% topology selection accuracy, <10% prediction error, and efficient inverse design—all within sub-second inference. The GNN forward model generalizes to unseen topologies with minimal fine-tuning, supporting broad practical deployment.
In future work, we aim to expand the topology library and support hierarchical macroblocks for scalable design beyond the cell level. We also plan to extend the layout-aware optimization with learned parasitic models and EM-informed constraints for more accurate post-layout estimation. Finally, integrating reinforcement learning or diffusion-based models for generative topology synthesis represents a promising step toward general-purpose analog design automation.
Acknowledgments
@@ Appendix A Qualitative Comparison with Prior Work
Method Topology Selection Parameter Inference Performance Prediction Layout Awareness Foundry Grade RF/ mm-Wave Public Dataset Public Code
@@
GCN-RL [10] ✔ (SPICE/Cadence)
Cao et al. [21] ✔ (SPICE/Cadence)
@@
RF/mm-Wave Is the method evaluated on at least one RF or mm-wave circuit type that reflects high-frequency design challenges?
@@ B.5 Voltage-Controlled Oscillators (VCOs)
Appendix C Graph-Based Circuit Representation
To enable GNN-based modeling of analog circuits, we represent each netlist as a directed multigraph where nodes correspond to electrical nets and edges encode circuit components such as transistors, capacitors, inductors, and voltage sources. Each edge is labeled with its component type and terminal role (e.g., gate, source, drain), and component-specific attributes are stored as edge features. For transistors, labels such as GS, DS, and DG denote source-to-gate, drain-to-source, and drain-to-gate connections, respectively.
Figure 13 illustrates two representative graph structures extracted from our dataset: an IFVCO and a ClassBPA. The visual encoding highlights the diversity of components and connectivity patterns across topologies. Edges corresponding to the same component type share a common color for visual consistency and semantic clarity. These structured graphs serve as the primary input to our GNN pipeline for performance prediction and inverse design.
(a) IFVCO
(b) ClassBPA
Figure 13: Graph representations of two analog circuit topologies from our dataset: (a) IFVCO and (b) ClassBPA. Nodes represent electrical nets, and colored edges denote circuit components such as transistors, capacitors, inductors, and sources. Each component type is visually distinguished by color and labeled with its name and terminal role (e.g., N2_GS, V0). These graphs serve as input to our GNN-based performance modeling and inverse design pipeline.
Appendix D Generalizing to Unseen Topologies via Fine-Tuning
To assess the generalization ability of our pretrained GNN, we evaluate it on the held-out RVCO topology, which was entirely excluded from the Stage 2 training, validation, and test splits (see Section 5). Notably, the RVCO training partition used here matches that of the Stage 1 experiments (Section 4), enabling consistent cross-stage evaluation.
Even in the zero-shot setting—where the model has never seen RVCO topologies—the pretrained GNN achieves a nontrivial mean relative error of 33.1%, highlighting its strong cross-topology generalization. Fine-tuning reduces this error to just 0.8%, demonstrating that the structural and parametric priors learned during pretraining are highly transferable. Table 11 reports detailed performance across five key metrics, confirming that the pretrained GNN can be rapidly adapted to novel circuit families with minimal supervision.
Metric DCP OscF TR OutP PN
Unit W GHz GHz dBm dBc/Hz
1.0 1.0 1.0 0.97 0.99
RMSE 0.725 0.244 0.022 0.098 0.792
MAE 0.576 0.19 0.018 0.078 0.477
Rel. Err. 0.85% 0.62% 1.4% 0.69% 0.56%

### v2

@@ 1 Introduction
  • We show that FALCON enables fast, reliable inverse design under layout and physical constraints, generating high-quality circuits in under one second per instance on CPU.
@@ 2 Related Work
To the best of our knowledge, FALCON is the first framework to unify topology selection, parameter inference, and layout-aware optimization in a single end-to-end pipeline, validated at scale using industrial-grade Cadence simulations for mm-wave analog circuits. A qualitative comparison with representative prior work is provided in Appendix A.
@@ 3.1 Dataset Overview
For each topology, expert-designed schematics were implemented in Cadence Virtuoso, and key design parameters were manually identified based on their functional relevance. Parameter ranges were specified by domain experts and systematically swept using Cadence ADE XL, enabling parallelized Spectre simulations across the design space. For each configuration, performance metrics—such as gain, bandwidth, and oscillation frequency—were extracted and recorded. Each datapoint therefore includes the full parameter vector, the corresponding Cadence netlist, and the simulated performance metrics. The resulting dataset comprises over one million datapoints, capturing a wide range of circuit behaviors and design trade-offs across diverse topologies. This large-scale, high-fidelity dataset forms the foundation for training and evaluating our inverse design pipeline. Detailed metric definitions and per-topology parameter ranges appear in Appendix B.
@@ 3.2 Graph-Based Circuit Representation
To support learning over such structured graphs, each edge is annotated with a rich set of attributes: (i) a categorical device type, specifying the component and connected terminal pair (e.g., NMOS drain–gate, resistor); (ii) numeric attributes, such as channel length or port resistance, fixed by the schematic; (iii) parametric attributes, defined symbolically in the netlist (e.g., W1, R3) and resolved numerically during preprocessing; (iv) one-hot categorical features, such as source type (DC, AC, or none); and (v) computational attributes, such as diffusion areas (Ad, As) derived from sizing. This rule-based graph construction generalizes across circuit families without task-specific customization. Graphs in the FALCON dataset range from 4–40 nodes and 7–70 edges, reflecting the variability of practical analog designs. Figure 2 shows two representative graph examples from our dataset—IFVCO and ClassBPA.
(a) IFVCO
(b) ClassBPA
Figure 2: Graph representations of two analog circuit topologies from our dataset: (a) IFVCO and (b) ClassBPA. Nodes represent electrical nets, and colored edges denote circuit components such as transistors, capacitors, inductors, and sources. Each component type is visually distinguished by color and labeled with its name and terminal role (e.g., N2_GS, V0). For transistors, labels such as GS, DS, and DG denote source-to-gate, drain-to-source, and drain-to-gate connections, respectively. These graphs serve as input to our GNN-based performance modeling and inverse design pipeline.
@@ 4 Stage 1: Performance-Driven Topology Selection
We assess classification performance using accuracy, balanced accuracy, macro precision, macro recall, macro F1, and micro F1 scores on the test set. As summarized in Table 1, the classifier achieves an overall accuracy of 99.57%, with macro F1 of 99.30% and balanced accuracy of 99.33%, demonstrating strong generalization across all 20 circuit topologies. Micro F1 (identical to accuracy in the multiclass setting) reaches 99.57%, while macro metrics—averaged equally across classes—highlight robustness to class imbalance. Seed-averaged results with 95% confidence intervals are provided in Appendix C. These trends are reinforced by the per-class accuracy plot in Figure 3(c), where most topologies reach 100% accuracy.
@@ 5 Stage 2: Generalizable Forward Modeling for Performance Prediction
where $m_{i}=1$ if the $i$-th metric is defined for the current sample, and 0 otherwise.
Model Architecture and Training. Each circuit is represented as an undirected multi-edge graph with voltage nets as nodes and circuit components as edges. All circuit parameters—both fixed and sweepable—are assigned to edges, along with categorical device types and one-hot encoded indicators. For each edge $e$, these attributes are concatenated to form a unified feature vector $x_{e}$. The feature set is consistent within each component type but varies across types (e.g., NMOS vs. inductor), reflecting the structure defined in Section 3.2.
@@
To account for component heterogeneity, we apply type-specific MLP encoders $\phi_{\text{enc}}^{(t_{e})}$ to the raw features $x_{e}$ of each edge $e$, producing edge embeddings $z_{e}=\phi_{\text{enc}}^{(t_{e})}(x_{e})$, where $t_{e}$ denotes the component type of edge $e$. Node features are initialized as scalar dummy values and mapped to a hidden dimension via a shared linear layer. We then apply a 4-layer edge-centric message-passing GNN with residual connections. At each layer $\ell$, every node $u$ receives messages from all incident edges $e\in\mathcal{E}_{u}$, where $\mathcal{E}_{u}$ denotes the set of edges connected to $u$. Messages are computed using a shared function $\phi_{\text{MSG}}$, which takes the hidden state of the neighboring node $\text{src}(e)$ and the associated edge embedding. Node states are then updated via a learnable function $\phi_{\text{UPD}}$, followed by a residual connection and nonlinearity:
$m_{u}^{(\ell)}=\sum_{e\in\mathcal{E}_{u}}\phi_{\text{MSG}}\left(h_{\text{src}(e)}^{(\ell)},z_{e}\right),\quad h_{u}^{(\ell+1)}=\text{ReLU}\!\left(\phi_{\text{UPD}}\!\left(m_{u}^{(\ell)}\right)+h_{u}^{(\ell)}\right)$
where $z_{e}\in\mathbb{R}^{d}$ is the learned embedding of edge $e$, and $h_{\text{src}(e)}^{(\ell)}\in\mathbb{R}^{d}$ is the hidden state of the node at the other end of edge $e$ from $u$ at layer $\ell$. The update function $\phi_{\text{UPD}}$ is a shared linear transformation applied to all nodes. After $L=4$ GNN layers, node embeddings are aggregated using global mean pooling to obtain a fixed-size graph representation:
$z_{\text{graph}}=\frac{1}{|V|}\sum_{u\in V}h_{u}^{(L)},$
where $V$ is the set of all nodes in the graph. The resulting vector is then passed through a fully connected MLP (output_mlp) to predict the 16-dimensional performance vector $\hat{y}\in\mathbb{R}^{16}$. An overview of this GNN-based forward prediction pipeline is shown in Figure 5.
@@
Evaluation. We evaluate the accuracy of the GNN forward model $f_{\theta}$ on a test set drawn from 19 of the 20 topologies. One topology—RVCO—is entirely excluded from training, validation, and test splits to assess generalization to unseen architectures. Prediction quality is measured using standard regression metrics: coefficient of determination ($\text{R}^{2}$), root mean squared error (RMSE), and mean absolute error (MAE), computed independently for each of the 16 performance metrics. We also report the mean relative error per metric, computed as the average across all test samples where each metric is defined. As summarized in Table 2, the model achieves high accuracy across all dimensions, with an average $\text{R}^{2}$ of 0.972.
@@
1.0 1.0 0.99 1.0 0.93 1.0 0.99 0.98 0.97 0.83 0.97 1.0 1.0 1.0 0.89 1.0
RMSE 0.27 0.101 0.536 0.833 1.515 0.21 0.534 0.972 0.723 0.293 0.91 0.095 0.226 0.143 2.536 0.071
MAE 0.198 0.072 0.208 0.188 0.554 0.12 0.2 0.376 0.184 0.097 0.238 0.066 0.163 0.105 1.159 0.046
Rel. Err. 11.2% 2.6% 19.0% 6.1% 11.4% 1.9% 4.5% 6.5% 0.6% 6.5% 4.6% 4.4% 4.6% 11.0% 1.3% 1.4%
To evaluate end-to-end prediction accuracy at the sample level, we compute the mean relative error per instance, defined as the average relative error across all valid (non-masked) performance metrics for each test sample. Figure 6 shows the distribution of this quantity across the test set (trimmed at the 95th percentile to reduce the impact of outliers). The distribution is sharply concentrated, indicating that most predictions closely match their corresponding target vectors. Without percentile trimming, the overall mean relative error across the full test set is 9.09%. Seed-averaged results with 95% confidence intervals are provided in Appendix C.
Generalizing to Unseen Topologies via Fine-Tuning. To assess the generalization ability of our pretrained GNN, we evaluate it on the held-out RVCO topology, which was entirely excluded from the Stage 2 training, validation, and test splits. Notably, the RVCO training partition used here matches that of the Stage 1 experiments (Section 4), enabling consistent cross-stage evaluation.
Even in the zero-shot setting—where the model has never seen RVCO topologies—the pretrained GNN achieves a nontrivial mean relative error of 30.4%, highlighting its strong cross-topology generalization. Fine-tuning reduces this error to just 0.9%, demonstrating that the structural and parametric priors learned during pretraining are highly transferable. Table 3 reports detailed performance across five key metrics of RVCO, confirming that the pretrained GNN can be rapidly adapted to novel circuit families with minimal supervision.
Metric DCP OscF TR OutP PN
Unit W GHz GHz dBm dBc/Hz
1.0 1.0 1.0 0.97 0.98
RMSE 0.643 0.324 0.026 0.099 0.953
MAE 0.508 0.256 0.02 0.077 0.619
Rel. Err. 0.75% 0.85% 1.63% 0.69% 0.73%
@@ 6 Stage 3: Layout-Aware Parameter Inference via Gradient Reasoning
where $\mathcal{L}_{\text{perf}}$ is the masked mean squared error (see Eqn 2) between predicted and target performance vectors, and $\mathcal{L}_{\text{layout}}$ is a normalized area penalty derived from analytical layout equations. To prioritize functionality, layout loss is softly gated by a sigmoid function:
@@
where $\sigma(\cdot)$ denotes the sigmoid function, and $\gamma$, $\tau$ are fixed scalars controlling the sharpness and center of the gating. This gating attenuates layout penalties when performance error exceeds a threshold $\tau$, encouraging the model to first achieve functionality before optimizing for layout compactness.
We set $\tau=0.05$, $\gamma=50$, and normalize layout area by $1\,\text{mm}^{2}$ to stabilize gradients. The layout weight $\lambda_{\text{area}}=0.02$ is chosen empirically to balance performance accuracy and physical realism without dominating the loss. This gated formulation supports manufacturable parameter recovery and reflects the broader paradigm of physics-informed learning [57]. Further discussion on user-defined objectives is provided in Appendix D.
@@
The layout model is deterministic and non-learned. It estimates area contributions from passive components—capacitors, inductors, and resistors—as these dominate total area and exhibit layout-sensitive behavior. Active devices (e.g., MOSFETs) are excluded since their geometries are fixed by the PDK and are negligible [62]. For a given parameter vector $x$, the total layout loss is computed as: $\mathcal{L}_{\text{layout}}(x)=\sum_{e\in\mathcal{E}_{\text{passive}}}A_{e}(x),$ where $\mathcal{E}_{\text{passive}}$ is the set of passive elements, and $A_{e}(x)$ is the area of the created layout for the passive component based on analytical physics-based equations. The area of element $e$ is estimated based on its 2D dimensions (e.g., $A=W\cdot L$ for resistors and capacitors). This area is normalized and used as a differentiable penalty in the optimization objective (see Eqn 3). Further implementation details are provided in Appendix E.
Gradient Reasoning Procedure. Starting from the initialized parameter vector $x_{0}$, we iteratively update parameters via gradient reasoning. At each step, the frozen forward model $f_{\theta}$ predicts the performance $\hat{y}=f_{\theta}(T,x)$, and the total loss $\mathcal{L}_{\text{total}}$ is evaluated. Gradients are backpropagated with respect to $x$, and updates are applied using the Adam optimizer. Optimization proceeds for a fixed number of steps, with early stopping triggered if the loss fails to improve over a predefined window.
@@ 7 Conclusion and Future Work
We presented FALCON, a modular framework for end-to-end analog and RF circuit design that unifies topology selection, performance prediction, and layout-aware parameter optimization. Trained on over one million Cadence-simulated mm-wave circuits, FALCON combines a lightweight MLP, a generalizable GNN, and differentiable gradient reasoning to synthesize circuits from specification to layout-constrained parameters. FALCON achieves >99% topology selection accuracy, <10% prediction error, and efficient inverse design—all within sub-second inference. In addition, the GNN forward model generalizes to unseen topologies with minimal fine-tuning, supporting broad practical deployment. Further discussion of training and inference efficiency, as well as practical limitations, is provided in Appendix F.
In future work, we aim to expand the topology library and support hierarchical macroblocks for scalable design beyond the cell level. We also plan to extend the dataset to cover multiple operating frequencies, enabling validation across diverse bands, and to enhance the layout-aware optimization with learned parasitic models, EM-informed constraints, and electromigration considerations for more accurate post-layout estimation. Finally, integrating reinforcement learning or diffusion-based models for generative topology synthesis represents a promising step toward general-purpose analog design automation.
Acknowledgments and Disclosure of Funding
@@ Appendix A Qualitative Comparison with Prior Work
Method
Topology
Selection
Parameter
Inference
Performance
Prediction
Layout
Awareness
Foundry
Grade
RF/
$m$-Wave
Public
Dataset
Public
Code
@@
GCN-RL [10] ✔ (SPICE/Cadence) ✘ (incomplete)
Cao et al. [21] ✔ (ADS/Cadence)
@@
RF/$m$-Wave Is the method evaluated on at least one RF or mm-wave circuit type that reflects high-frequency design challenges?
@@ B.5 Voltage-Controlled Oscillators (VCOs)
Appendix C Robustness Across Random Seeds
To evaluate the robustness of our models to random initialization and data shuffling, we repeated experiments using five distinct random seeds: {42, 123, 777, 2023, 3407}. Reporting across multiple seeds is important for ensuring that observed results are not artifacts of a specific initialization or training trajectory, but rather reflect the stable behavior of the method. For each metric, we compute the mean and 95% confidence interval across seeds, reporting results in the form $\mu\pm\Delta$.
Table 12: Topology selection performance with mean scores and 95% confidence intervals across five random seeds.
Metric Mean $\pm$ 95% CI (%)
Accuracy 99.57 $\pm$ 0.01
Balanced Accuracy 99.34 $\pm$ 0.02
Macro Precision 99.27 $\pm$ 0.01
Macro Recall 99.34 $\pm$ 0.02
Macro F1 99.30 $\pm$ 0.01
Micro F1 99.57 $\pm$ 0.01
For the MLP topology selection model, results are highly stable across random seeds. The accuracy reaches $99.57\pm 0.01\%$ with balanced accuracy at $99.34\pm 0.02\%$, while both macro and micro F1 scores exceed $99.3\%$ with confidence intervals no larger than $\pm 0.02$. These narrow intervals indicate that the MLP’s performance is effectively invariant to random initialization, underscoring its robustness and reliability in the topology selection stage of the pipeline (Section 4).
For the GNN-based forward performance prediction model, the overall mean relative error across all metrics is $9.14\pm 0.38\%$ (95% CI). Individual performance predictions, including DC power consumption, gain, bandwidth, and oscillation frequency, exhibit narrow confidence intervals—for example, noise figure achieves $4.48\pm 0.07\%$ error and oscillation frequency $0.65\pm 0.03\%$. These results indicate that the GNN achieves consistently accurate predictions across diverse circuit characteristics. The tight confidence intervals further demonstrate that the model’s performance is robust to random initialization, underscoring its reliability as a generalizable forward predictor within the pipeline (Section 5). The full seed-dependent results for both models are provided in Tables 12 and 13.
Table 13: Prediction accuracy of the forward GNN with mean scores and 95% confidence intervals across five random seeds.
Rel. Err.
$\pm$ 95% CI (%)
11.64
$\pm$ 1.06
3.10
$\pm$ 0.42
18.46
$\pm$ 0.36
5.25
$\pm$ 0.44
11.49
$\pm$ 0.09
1.94
$\pm$ 0.15
4.48
$\pm$ 0.07
6.28
$\pm$ 0.43
0.65
$\pm$ 0.03
6.55
$\pm$ 0.04
4.86
$\pm$ 0.59
4.31
$\pm$ 0.24
4.51
$\pm$ 0.14
11.58
$\pm$ 1.72
1.34
$\pm$ 0.02
1.71
$\pm$ 0.29
Appendix D User-Defined Loss Functions for Gradient Reasoning
Stage 3 of FALCON employs gradient reasoning with the forward GNN fixed, enabling the optimization objective to be redefined without retraining or fine-tuning the predictive model. This design allows users to flexibly adapt the loss function to capture specific trade-offs or constraints. We illustrate this flexibility with two examples.
Weighted Performance Loss. Rather than treating all performance metrics equally, users can specify weights $\alpha_{i}$ for each target metric:
$\mathcal{L}_{\text{perf-weighted}}=\frac{1}{\sum_{i}m_{i}\alpha_{i}}\sum_{i=1}^{d}m_{i}\alpha_{i}\,(\hat{y}_{i}-y_{i}^{\text{target}})^{2},$
where larger $\alpha_{i}$ prioritize certain specifications (e.g., gain or noise figure). Here, $m_{i}=1$ if the $i$-th metric is defined for the current sample, and 0 otherwise.
Interval-Constrained Performance Loss. Users may also define acceptable ranges for metrics rather than fixed targets. Given optional lower and/or upper bounds $y_{i}^{\text{lower}},y_{i}^{\text{upper}}$, the interval penalty is:
$\mathcal{L}_{\text{perf-interval}}=\frac{1}{\sum_{i}m_{i}}\sum_{i=1}^{d}m_{i}\Big[\mathbbm{1}_{\{y_{i}^{\text{upper}}\,\text{defined}\}}\max(0,\hat{y}_{i}-y_{i}^{\text{upper}})+\mathbbm{1}_{\{y_{i}^{\text{lower}}\,\text{defined}\}}\max(0,y_{i}^{\text{lower}}-\hat{y}_{i})\Big],$
where the indicator $\mathbbm{1}_{\{\cdot\}}$ indicates whether the corresponding bound is specified. This formulation naturally handles the cases where only an upper bound, only a lower bound, or both bounds are provided. As above, $m_{i}=1$ if the $i$-th metric is defined for the current sample, and 0 otherwise.
General Extensibility. More generally, the total loss in Eqn. 3 can be replaced with any user-defined formulation, allowing both $\mathcal{L}_{\text{perf}}$ and $\mathcal{L}_{\text{layout}}$ to be substituted with customized objectives. Additional physical constraints, multi-objective trade-offs, or alternative layout penalties can be incorporated with only a few lines of code. This extensibility underscores the flexibility of FALCON and enables the framework to adapt to diverse design objectives.